OpenStudy (anonymous):

. sq. rt of ( 2x - 1) = x

OpenStudy (anonymous):

Make the substitution u = sqrt(1 + 2x). Then du = (1 / sqrt(1 + 2x)) dx. Solve the first equation for x: u = sqrt(1 + 2x) u^2 = 1 + 2x u^2 - 1 = 2x (u^2 - 1) / 2 = x. Now you can plug in for x and dx in your integral: ∫ (x / sqrt(1 + 2x)) dx = ∫ [(u^2 - 1) / 2] du. This can be integrated with the power rule: ∫ [(u^2 - 1) / 2] du = (1/2) ∫ (u^2 - 1) du = (1/2)(u^3 / 3 - u) + C = u^3 / 6 - u / 2 + C. Now substitute back in for u: u^3 / 6 - u / 2 + C = (sqrt(1 + 2x))^3 / 6 - (sqrt(1 + 2x)) / 2 + C. = [(1 + 2x)^(3/2)] / 6 - (sqrt(1 + 2x)) / 2 + C.

OpenStudy (anonymous):

i have done integration !!oops!!

OpenStudy (anonymous):

so thats the wrong way ?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

?

OpenStudy (radar):

\[\sqrt{2x-1}=x\]Square both sides getting\[2x-1=x ^{2}\] Transpose by subtracting (2x-1) from both sides.\[x ^{2}-2x+1=0\]

OpenStudy (radar):

Factor getting (x-1)(x-1)=0 x=1 (double root)

OpenStudy (radar):

check by substituting in original.\[\sqrt{2(1)-1}=1\]\[\sqrt{1}=1\]